OpenStudy (anonymous):

Write 8/13 and 10/17 as the sum of three or more unit fractions. There has to be a better way to do this than trial and error...

OpenStudy (anonymous):

By unit fractions you mean fractions of this type ... (1/a) where a is an integer?

OpenStudy (anonymous):

Lol... best I can think of is (8/13) = (1/13) + (1/13) ... eight times (10/17) = (1/17) + (1/17) ... ten times Meets problem requirements. As 13 and 17 are prime numbers, this is kind of hard to do otherwise... no ideas.